Suppose that for a particular economy and period, investment was equal to 200, government expenditure was equal to 100, net taxes were fixed at 150, and consumption (C) was given by the consumption function C = 20 + 0.6YD where YD is disposable income and Y is GDP. a. What is the level of equilibrium income (Y) ? b. What is the value of the government expenditure multiplier (Y>G) ? Of the tax multiplier (Y>T)? c. Suppose the investment declined by 100 units to a level of 100. What will be the new level of equilibrium income?
1
Expert's answer
2019-10-03T09:24:01-0400

a. Y = C + I + G + (X – M)= 20 + 0.6(Y-150)+200+100 = 20+0.6Y-90+200+100

Y-0.6Y= 230

Y= "\frac{230}{0.4}" = 575


b. government expenditure multiplier= "\frac{575}{100}" = 5.75


 tax multiplier ="\frac{575}{150}" = 3.833


c. Y= "\frac{130}{0.4}" = 325